Unpolarized light with an intensity of 120 W falls on a polarizer and analyser. The angle between the transmission axes of the polarizer and analyser is 72°. What is the intensity of the light leaving the analyser, reaching the photocell?

Question 3 options:


A) 

No option is correct


B) 

11 W


C) 

5.7 W


D) 

60 W


E) 

19 W


1
Expert's answer
2021-11-11T17:09:16-0500

"I=\frac{1}{2}I_0\cos^2\alpha=\frac{1}{2}\cdot120\cdot\cos^272°=5.7\ (W)"


C) 5.7 W . Answer